Buhle wants to save R330 284,00 within the next four years. She makes equal monthly deposits at the end of each month starting after one month and her savings accumulate interest at 12,5% per annum, compounded monthly. The size of each of the monthly deposits is


1
Expert's answer
2021-04-26T05:41:02-0400

"FV A=A×(\frac{(1+\frac{r}{m)}^{m×n}-1}{\frac{r}{m}})"

future value(Fv) = R330 284,00

Interest rate(r)=12.5%

m=12 (monthly compounding)

Time(n )= 4 years


let the size of monthly payment = A,

"330 284,00=A×\frac{(1+\frac{0.125}{12})^{12×4}-1}{\frac{0.125}{12}}"


"33028400=A×(\frac{1.0104167^{48}-1}{0.0104167})"


"33028400=A\u00d761.87"


"A=\frac{330 284,00}{61.87}"


"533849"